# Debt Behind the AI Boom: A Large-Scale Empirical Study of AI-Generated Code in the Wild

### 30 Mar 2026 v1

@@ Abstract.
AI coding assistants are now widely used in software development. Software developers increasingly integrate AI-generated code into their codebases to improve productivity. Prior studies have shown that AI-generated code may contain code quality issues under controlled settings. However, we still know little about the real-world impact of AI-generated code on software quality and maintenance after it is introduced into production repositories. In other words, it remains unclear whether such issues are quickly fixed or persist and accumulate over time as technical debt. In this paper, we conduct a large-scale empirical study on the technical debt introduced by AI coding assistants in the wild. To achieve that, we built a dataset of 304,362 verified AI-authored commits from 6,275 GitHub repositories, covering five widely used AI coding assistants. For each commit, we run static analysis before and after the change to precisely attribute which code smells, bugs, and security issues the AI introduced. We then track each introduced issue from the introducing commit to the latest repository revision to study its lifecycle. Our results show that we identified 484,606 distinct issues, and that code smells are by far the most common type, accounting for 89.1% of all issues. We also find that more than 15% of commits from every AI coding assistant introduce at least one issue, although the rates vary across tools. More importantly, 24.2% of tracked AI-introduced issues still survive at the latest revision of the repository. These findings show that AI-generated code can introduce long-term maintenance costs into real software projects and highlight the need for stronger quality assurance in AI-assisted development.
AI coding assistants, technical debt, code quality, software maintenance, empirical study
copyright: none
1. Introduction
@@
Through AI coding assistants (e.g., Cursor, Claude Code), software developers can now describe what they want in natural language and get working code back in seconds. They significantly improve development productivity. Thus, AI is becoming standard equipment for modern software developers (GitHub, 2024). According to the 2025 Stack Overflow Developer Survey, 84% of professional developers are using or are planning to use AI tools within their development processes (Stack Overflow, 2025). The AI-generated code is also widely used in real-world software projects. For example, both Google and Microsoft disclosed in 2025 that AI now writes over 20% of their new code (Novet, 2025; Robison, 2024). Similarly, GitHub reported that more than 1.1 million public repositories used AI coding tools between 2024 and 2025  (GitHub, 2024). Overall, AI-generated code has graduated from experiment to production reality, and it is everywhere.
@@
Although AI coding assistants have proven effective at generating functional programs, many previous research studies have revealed a range of quality concerns in AI-generated code. Recent studies have shown that AI-generated code suffers from functional bugs, runtime errors, and systemic maintainability issues (Liu et al., 2024; Siddiq et al., 2024). Also, the code produced by AI tools poses security risks (Pearce et al., 2025; Perry et al., 2023; Liu et al., 2025). Pearce et al. (Pearce et al., 2025) found that about 40% of AI-generated code in security-sensitive contexts contains critical vulnerabilities. However, recent research (Perry et al., 2023; Sabouri et al., 2025) has found that developers tend to place excessive trust in the quality of AI-generated code, blindly accepting the code without proper validation. As a result, these unverified code snippets are merged into production codebases. Over time, this can cause a considerable accumulation of technical debt, which can be costly and time-consuming to address (Harding and Kloster, 2024; Moreschini et al., 2026).
@@
Recognizing these risks, recent empirical studies have started to investigate AI-generated code in real-world repositories. These studies cover a range of practical concerns, such as security weaknesses (Fu et al., 2025; Wang et al., 2025), project-level development velocity (He et al., 2025), pull request acceptance rates (Watanabe et al., 2025), and code redundancy (Huang et al., 2026). For example, He et al. (He et al., 2025) found that Cursor adoption in 807 GitHub repositories led to a transient velocity boost but persistent increases in code complexity. However, existing studies still have several limitations. First, most studies focus on a single tool or a narrow set of tools, which limits the generalizability of their findings. Second, many studies rely on proxy-based attribution (e.g., configuration files, self-disclosure signals, or classifiers), rather than directly tracking which code lines or files are AI-generated. Third, they mostly evaluate code at a single point in time, failing to capture the long-term lifecycle of the code. Thus, it remains unknown how AI-generated code actually ages in production. We still do not know whether the technical debt it introduces persists, gets refactored, or silently accumulates over time.
@@
To bridge this gap, we conduct a large-scale empirical study to investigate the lifecycle of AI-introduced technical debt in the wild. Our approach consists of three steps. First, we build a large dataset of verified AI-authored commits. Instead of relying on classifiers or proxy signals, we use explicit Git metadata to identify commits generated by AI coding tools across thousands of GitHub repositories. Second, we perform a commit-level quality analysis. For each AI-authored commit, we run static analysis tools on the source code immediately before and after the change. This allows us to precisely identify which code smells, bugs, and security issues the AI introduced or fixed. Third, we conduct a debt lifecycle analysis. We track each introduced issue to the latest repository revision (HEAD) to determine whether it still survives or has been resolved. Through this design, our study provides the first comprehensive view of how AI-generated code ages in real-world software.
@@
  • Perform commit-level differential analysis to precisely attribute code smells, bugs, and security issues to individual AI-authored commits.
@@
  • Track the lifecycle of AI-introduced technical debt from the moment of introduction to the latest repository revision, revealing how it persists, evolves, or gets resolved over time.
@@
Open Science. To support the open science initiative, we publish the studies dataset and a replication package, which is publicly available in GitHub. [^1] https://github.com/xxxx/xxxx
@@
Paper Organization. Section 2 presents the background and motivation. Section 3 describes our approach. Section 4 presents the experimental setup. Section 5 presents the results. Section 6 discusses the findings and their implications. Section 7 presents the related work. Section 8 discloses the threats to validity. Section 9 draws the conclusions.
2. Background
2.1. AI Coding Assistants
@@
Modern AI coding assistants (e.g., Cursor, Claude Code, GitHub Copilot) are now deeply embedded in software development workflows. They help software developers write, modify, explain, test, or debug code using natural-language instructions and code context. Advanced agentic tools can now process entire functions, files, or even entire repositories to autonomously create pull requests with minimal human intervention. Driven by these improved capabilities, AI-generated code is entering production codebases at unprecedented speed and scale. According to GitHub, over 1.1 million public repositories adopted AI coding tools between 2024 and 2025 (GitHub, 2024). As shown in Figure 1, in Anthropic’s claudes-c-compiler repository (Anthropic, 2026), Claude appears as the top contributor, with 3,957 commits and nearly 500K lines of code added within just a few weeks. At this volume and speed, it is unlikely that all AI-generated code receives a thorough human review. This makes it increasingly important to understand the long-term implications of AI-generated code on software quality and maintenance.
Figure 1. Contributor statistics for the Anthropic claudes-c-compiler repository.
@@
In controlled (lab) settings (Pearce et al., 2025; Liu et al., 2024), the provenance of code is usually clear since researchers can generate the code directly. However, it is different in real-world codebases. The reason is that AI-generated code and human-written code are often mixed together. This makes it harder to observe the long-term impact of AI-generated code after it is merged into production codebases.
$2.$ Technical Debt
@@
Technical debt refers to design or implementation choices that prioritize short-term speed over long-term quality (Cunningham, 1992). These shortcuts may help in the short term, but they increase the future cost of maintaining and evolving the software (Avgeriou et al., 2016; Li et al., 2015). Its costs can accumulate over time if not addressed. This concern becomes more important as AI coding assistants are widely adopted. AI coding assistants help developers work faster and produce more code. Previous studies have shown that AI-generated code contains code smells, maintainability issues, and security vulnerabilities (Liu et al., 2024; Pearce et al., 2025; Fu et al., 2025). When such issues are accepted into production repositories, they can accumulate as technical debt in the codebase.
seagullz4/hysteria2 (seagullz4, 2025c), hysteria2.py:L77, e277daf $\Downarrow$ seagullz4/hysteria2, hysteria2.py:L77 d9e392d
@@
Figure 2. Command injection risk introduced by GitHub Copilot in hysteria2 (1.7K stars) and its subsequent fix.
realsenseai/librealsense (RealSense, 2025), test-fps-performance.py:L168, 5535b8a $\Downarrow$ realsenseai/librealsense, test-fps-performance.py:L37, 14026c8
@@
Figure 3. Undefined variables introduced by GitHub Copilot in librealsense (8.6K stars), causing a runtime error. Fixed three weeks later.
2.3. Motivation
@@
The technical debt introduced by AI coding assistants is not just a theoretical risk. Such issues are becoming increasingly common in real-world codebases. Figure 2 and 3 show two examples. In Figure 2, a Copilot-authored commit introduced a shell=True subprocess call in hysteria2.py (seagullz4, 2025b). This pattern increases security risk by allowing command injection if user input is involved. A human developer later fixed it, noting in the commit message: “Improve code security by removing shell=True from subprocess calls” (seagullz4, 2025a). Figure 3 shows a bug from Intel’s librealsense (RealSense, 2025). A Copilot-authored commit replaced a value with a named constant but never defined it (Intel RealSense, 2025b). The buggy code remained in the repository for over three weeks before the maintainer added the missing definition (Intel RealSense, 2025a).
@@
These two examples highlight the motivation of our study. AI coding assistants can generate functional code, but they may introduce quality issues into production codebases. Developers also tend to over-trust and accept AI suggestions without thorough review (Perry et al., 2023; Sabouri et al., 2025). These issues may be fixed later, or they may persist for a long time, or even create long-term maintenance challenges. Recent studies (Nguyen and Nadi, 2022; Liu et al., 2024; Fu et al., 2025; Siddiq et al., 2024; Wang et al., 2025; He et al., 2025) have examined the AI-generated code, but they have several limitations. However, most prior studies focus on a single tool, a small set of tasks, or controlled settings. For example, Watanabe et al. (Watanabe et al., 2025) measured the initial acceptance rate of AI-authored code in a single repository. We still know little about the long-term implications of AI-generated code on software quality and maintenance. Thus, we study the technical debt introduced by AI coding assistants in the wild: how often it is introduced, what kinds of issues appear, and whether those issues still remain in the codebase over time.
3. Approach
Figure 4. Overview of our approach.
@@ 3. Approach
Figure 4 provides an overview of our approach. We first collect AI-authored commits from GitHub repositories at scale (Section 3.1). We then analyze each AI-authored commit at the code level to determine which quality issues it introduced or fixed (Section 3.2). Finally, we track the lifecycle of both the issues and the code itself to determine whether AI-introduced debt persists or gets resolved over time (Section 3.3).
3.1. Data Collection
@@
Repository Discovery.  We use the GitHub Archive dataset, which records all public events (i.e., PushEvent) on GitHub, to identify repositories with potential AI-authored code. We scan all PushEvent records from January 2024 to October 2025, focusing on repositories with recent development activity. Using Google BigQuery, we extract four metadata fields from each event: actor login, author name, author email, and commit message. We then match these fields against our curated AI-attribution patterns. Only repositories with at least one matching event are retained. It is worth mentioning that GitHub changed the Events API on October 7, 2025, and removed commit-level summaries from PushEvent payloads (GitHub, 2025). Thus, for the post-October-2025 data, we supplement BigQuery-based discovery with GitHub API queries over top repositories and then apply the same attribution rules through full-history repository scanning (described in the next section). These steps help us ensure that active and popular repositories are not missed due to the API change.
AI Attribution Rules.  We build attribution rules for widely adopted AI coding tools (e.g., Cursor, GitHub Copilot, Claude Code) identified in the 2025 Stack Overflow Developer Survey (Stack Overflow, 2025). We identify AI-authored commits using explicit signals in Git metadata. Our approach covers AI-authored commits only when the use of an AI coding tool leaves explicit traces in Git metadata The rules are based on four sources of evidence: (1)  actor logins (e.g., copilot-swe-agent[bot]), (2) author emails (e.g., noreply@anthropic.com), (3) author names (e.g., Cursor Agent), and (4) Co-authored-by trailers in commit messages. These signals identify whether and which AI coding tool was involved in a commit. In total, 29 AI coding tools left identifiable traces in the repositories we collected. The full list of tools and detection rules is included in our replication package.
@@
Filtering.  To focus on established open-source projects, we filter out repositories that do not meet our study criteria. We keep only repositories with at least 100 GitHub stars. We remove forks with fewer than 100 stars. We also require at least one confirmed AI-authored commit. Our downstream analysis is restricted to production Python, JavaScript, and TypeScript source files, excluding repositories or commits that do not contain analyzable source code in these languages. After filtering, our dataset contains 6,275 repositories with 304.4k total AI-authored commits.
3.2. Commit-Level Quality Analysis
To understand the impact of AI coding tools, we avoid evaluating a repository snapshot at a single point in time. For each AI-authored commit $c$, we analyze the source code at its parent revision (before) and at the commit itself (after) to determine which quality issues it introduced or fixed. Our analysis focuses on production source files written in Python, JavaScript, and TypeScript. We exclude files that are unlikely to reflect production code quality (e.g., tests, documentation, configuration files, auto-built artifacts, and vendored dependencies). Files are classified based on their paths and naming patterns (e.g., files under test/ or __tests__/ directories, or matching *_test.py). The full classification rules are included in our replication package.
Static Analysis.  For each AI-authored commit $c$ that modifies a source file $f$, we check out two versions of $f$ (i.e., the version before $c$ (at its parent commit) and the version after $c$). We run the same static analysis toolchain on both versions to identify code quality and security issues. For Python, we use Pylint and Bandit. For JavaScript and TypeScript, we use ESLint and njsscan. For each detected issue, we record its rule identifier, line number, severity, and message. This produces two issue sets II: the issue set before the commit, denoted by $I_{f}^{-}$, and the issue set after the commit, denoted by $I_{f}^{+}$.

### 26 Apr 2026 v2

@@ Abstract
AI coding assistants are now widely used in software development. Software developers increasingly integrate AI-generated code into their codebases to improve productivity. Prior studies have shown that AI-generated code may contain code quality issues under controlled settings. However, we still know little about the real-world impact of AI-generated code on software quality and maintenance after it is introduced into production repositories. In other words, it remains unclear whether such issues are quickly fixed or persist and accumulate over time as technical debt. In this paper, we conduct a large-scale empirical study on the technical debt introduced by AI coding assistants in the wild. To achieve that, we built a dataset of 302.6k verified AI-authored commits from 6,299 GitHub repositories, covering five widely used AI coding assistants. For each commit, we run static analysis before and after the change to precisely attribute which code smells, correctness issues, and security issues the AI introduced. We then track each introduced issue from the introducing commit to the latest repository revision to study its lifecycle. Our results show that we identified 484,366 distinct issues, and that code smells are by far the most common type, accounting for 89.3% of all issues. We also find that more than 15% of commits from every AI coding assistant introduce at least one issue, although the rates vary across tools. More importantly, 22.7% of tracked AI-introduced issues still survive at the latest version of the repository. These findings show that AI-generated code can introduce long-term maintenance costs into real software projects and highlight the need for stronger quality assurance in AI-assisted development.
I Introduction
@@
Through AI coding assistants (e.g., Cursor, Claude Code), software developers can now describe what they want in natural language and get working code back in seconds. They significantly improve development productivity. Thus, AI is becoming standard equipment for modern software developers [23]. According to the 2025 Stack Overflow Developer Survey, 84% of professional developers are using or are planning to use AI coding tools within their development processes [63]. The AI-generated code is also widely used in real-world software projects. For example, both Google and Microsoft disclosed in 2025 that AI now writes over 20% of their new code [45, 55]. Similarly, GitHub reported that more than 1.1 million public repositories used AI coding tools between 2024 and 2025  [23]. Overall, AI-generated code has graduated from experiment to production reality, and it is everywhere.
@@
Although AI coding assistants have proven effective at generating functional programs, many previous research studies have revealed a range of quality concerns in AI-generated code. Recent studies have shown that AI-generated code suffers from functional bugs, runtime errors, and systemic maintainability issues [35, 62]. Also, the code produced by AI coding tools poses security risks [47, 49, 36]. Pearce et al. [47] found that about 40% of AI-generated code in security-sensitive contexts contains critical vulnerabilities. However, recent research [49, 56] has found that developers tend to place excessive trust in the quality of AI-generated code, blindly accepting the code without proper validation. As a result, these unverified code snippets are merged into production codebases. Over time, this can cause a considerable accumulation of technical debt, which can be costly and time-consuming to address [27, 41].
@@
Recognizing these risks, recent empirical studies have started to investigate AI-generated code in real-world repositories. These studies cover a range of practical concerns, such as security weaknesses [20, 69], project-level development velocity [28], pull request acceptance rates [70], and code redundancy [29]. For example, He et al. [28] found that Cursor adoption in 807 GitHub repositories led to a transient velocity boost but persistent increases in code complexity. However, existing studies still have several limitations. First, most studies focus on a single tool or a narrow set of tools, which limits the generalizability of their findings. Second, many studies do not identify AI-generated code at the commit level. Instead, they rely on project-level adoption signals, such as the presence of AI tool configuration files in a repository (e.g., .cursorrules or .claude/). These signals suggest possible AI use, but do not directly show which commits or files were generated by AI. Third, they mostly evaluate code at a single point in time, failing to capture the long-term lifecycle of the code. Finally, in real-world repositories, AI-assisted code and human-written code are often mixed together, and AI usage may leave no reliable trace. Thus, it remains unknown how AI-generated code actually ages in production. We still do not know whether the technical debt it introduces persists, gets refactored, or silently accumulates over time.
@@
To bridge this gap, we conduct a large-scale empirical study to investigate the lifecycle of AI-introduced technical debt in the wild. Our approach consists of three steps. First, we build a large dataset of verified AI-authored commits from five popular coding assistants (i.e., GitHub Copilot, Claude, Cursor, Gemini, and Devin) across over 6,000 GitHub repositories. We use explicit Git metadata to identify commits generated by AI coding tools across thousands of GitHub repositories. This design is analogous in spirit to research on self-admitted technical debt (SATD), which studies the subset of technical debt that developers explicitly document rather than the full universe of debt [50, 11]. Similarly, we focus on the subset of AI-assisted contributions whose AI involvement is explicitly visible in Git metadata. Although this does not cover all AI-assisted changes, it provides a more reliable basis for attribution and large-scale empirical analysis. Second, we perform a commit-level quality analysis. For each AI-authored commit, we run static analysis tools on the source code immediately before and after the change. This allows us to precisely identify which code smells, correctness issues, and security issues the AI introduced or fixed. Third, we conduct a debt lifecycle analysis. We track each introduced issue to the latest repository revision (HEAD) to determine whether it still survives or has been resolved. Through this design, our study provides the first comprehensive view of how AI-generated code ages in real-world software.
@@
  • Perform commit-level differential analysis to precisely attribute code smells, correctness issues, and security issues to individual AI-authored commits.
@@
  • Track AI-introduced issues from the introducing commit to the latest repository revision, revealing whether it persists or gets resolved over time.
@@
Open Science. To support the open science initiative, we publish the studies dataset and a replication package, which is publicly available in GitHub. [^1] https://github.com/yueyueL/tech-debt-ai-coding
@@
Paper Organization. Section II presents the background and motivation. Section III describes our approach. Section IV presents the experimental setup. Section V presents the results. Section VI discusses the findings and their implications. Section VII presents the related work. Section VIII discloses the threats to validity. Section IX draws the conclusions.
II Background
II-A AI Coding Assistants
@@
Modern AI coding assistants (e.g., Cursor, Claude Code, GitHub Copilot) are now deeply embedded in software development workflows. They help software developers write, modify, explain, test, or debug code using natural-language instructions and code context. Advanced agentic tools can now process whole functions, files, or even repositories to autonomously create pull requests with minimal human intervention. Driven by these improved capabilities, AI-generated code is entering production codebases at unprecedented speed and scale. According to GitHub, over 1.1 million public repositories adopted AI coding tools between 2024 and 2025 [23]. As shown in Figure 1, in Anthropic’s claudes-c-compiler repository [2], Claude appears as the top contributor, with 3,957 commits and nearly 500K lines of code added within just a few weeks. At this volume and speed, it is unlikely that all AI-generated code receives a thorough human review. This makes it increasingly important to understand the long-term implications of AI-generated code on software quality and maintenance.
Figure 1: Contributor statistics for the Anthropic claudes-c-compiler repository.
@@
In controlled (lab) settings [47, 35], the provenance of code is usually clear since researchers can generate the code directly. However, it is different in real-world codebases. AI-generated code and human-written code are often interleaved during development, and AI usage is not always explicitly recorded in the repository history. This makes it harder to attribute code changes reliably and to observe the long-term impact of AI-generated code after it is merged into production codebases.
II-B Technical Debt
@@
Technical debt refers to design or implementation choices that prioritize short-term speed over long-term quality [9]. These shortcuts may help in the short term, but they increase the future cost of maintaining and evolving the software [6, 33]. Its costs can accumulate over time if not addressed. This concern becomes more important as AI coding assistants are widely adopted. AI coding assistants help developers work faster and produce more code. Previous studies have shown that AI-generated code contains code smells, correctness issues, and security vulnerabilities [35, 47, 20]. In this study, we focus on these three categories as code-level technical debt: code smells that reduce maintainability, correctness issues that affect program behavior, and security issues that may expose systems to risk. When such issues are accepted into production repositories, they can accumulate as technical debt in the codebase.
hysteria2 [60], hysteria2.py:L77, e277daf $\Downarrow$ hysteria2 [60], hysteria2.py:L77 d9e392d
@@
Figure 2: Command injection risk introduced by GitHub Copilot in hysteria2 (1.7K stars) [60] and later removed in a Copilot-assisted fix commit [58].
librealsense [53], test-fps-performance.py:L168, 5535b8a $\Downarrow$ librealsense [53], test-fps-performance.py:L37, 14026c8
@@
Figure 3: Undefined variables introduced by GitHub Copilot in librealsense (8.6K stars) [53], causing a runtime error. A human maintainer committed a fix three weeks later [30].
II-C Motivation
@@
The technical debt introduced by AI coding assistants is not just a theoretical risk. Such issues are becoming increasingly common in real-world codebases. Figure 2 and 3 show two examples. In Figure 2, a GitHub Copilot-authored commit introduced a shell=True subprocess call in hysteria2.py [59]. This pattern increases security risk by allowing command injection if user input is involved. A human developer later fixed it, noting in the commit message: “Improve code security by removing shell=True from subprocess calls” [58]. Figure 3 shows a bug from Intel’s librealsense [53]. In another case, GitHub Copilot replaced a literal value with a named constant, but never defined the constant [31]. The buggy code remained in the repository for over three weeks before the maintainer committed a fix adding the missing definition [30].
@@
These two examples highlight the motivation of our study. AI coding assistants can generate functional code, but they may introduce quality issues into production codebases. Developers also tend to over-trust and accept AI suggestions without thorough review [49, 56]. These issues may be fixed later, or they may persist for a long time, or even create long-term maintenance challenges. Recent studies [44, 35, 20, 62, 69, 28] have examined the AI-generated code, but they have several limitations. Most prior studies focus on a single tool, a small set of tasks, or controlled settings. For example, Watanabe et al. [70] measured the initial acceptance rate of AI-authored code in a single repository. We still know little about the long-term implications of AI-generated code on software quality and maintenance. Thus, we study the technical debt introduced by various AI coding assistants in the wild: how often it is introduced, what kinds of issues appear, and whether those issues still remain in the codebase over time.
III Approach
Figure 4: Overview of our approach.
@@ III Approach
Figure 4 provides an overview of our approach. We first collect AI-authored commits from GitHub repositories at scale (Section III-A). We then analyze each AI-authored commit at the code level to determine which quality issues it introduced or fixed (Section III-B). Finally, we track the lifecycle of both the issues and the code itself to determine whether AI-introduced debt persists or gets resolved over time (Section III-C).
III-A Data Collection
@@
Repository Discovery.  We use the GitHub Archive dataset [21], which records public GitHub events (e.g., PushEvent), to identify repositories with potential AI-authored code. The dataset is publicly available through Google BigQuery [22], which allows large-scale querying over historical events. We scan PushEvent records from January 2024 to October 2025, focusing on repositories with recent development activity. From each event, we extract four metadata fields: actor login, author name, author email, and commit message. We then match these fields against our curated AI-attribution rules (described in the next section) to identify potential AI-authored commits. Only repositories with at least one matching event are retained. To improve coverage of active and popular repositories, we also query the GitHub REST API [25] for top-starred repositories and apply the same attribution rules through full-history repository scanning. This step helps us ensure that repositories with substantial AI-authored activity are not missed during the discovery stage.
AI Attribution Rules.  We build attribution rules for widely adopted AI coding tools (e.g., Cursor, GitHub Copilot, Claude Code) identified in the 2025 Stack Overflow Developer Survey [63]. We identify AI-authored commits using explicit signals in Git metadata. Our approach covers AI-authored commits only when the use of an AI coding tool leaves explicit traces in Git metadata. The rules are based on four sources of evidence: (1)  actor logins (e.g., copilot-swe-agent[bot]), (2) author emails (e.g., noreply@anthropic.com), (3) author names (e.g., Cursor Agent), and (4) Co-authored-by trailers in commit messages. These rules rely on explicit machine-readable signals in Git metadata. To finalize the rule list, two authors manually reviewed candidate patterns and verified that they provided reliable evidence of AI coding tool involvement. In total, 29 AI coding tools left identifiable traces in the repositories we collected. The full list of tools and detection rules is included in our replication package.
@@
Filtering.  To focus on established open-source projects, we filter out repositories that do not meet our study criteria. We keep only repositories with at least 100 GitHub stars. We also require at least one confirmed AI-authored commit. Our downstream analysis is restricted to production Python, JavaScript, and TypeScript source files, since these are among the most widely used programming languages [63] and are well supported by static analysis tools. We therefore exclude repositories that do not contain any source files in these languages. In total, the discovery stage identified 587,118 candidate repositories. After applying the star threshold, 12,770 repositories remained. After full-history scanning and language filtering, we obtained 6,699 repositories with confirmed AI-authored commits.
III-B Commit-Level Quality Analysis
To understand the impact of AI coding tools, we avoid evaluating a repository snapshot at a single point in time. For each AI-authored commit $c$, we analyze two versions of the source code: the version at $c$’s parent revision (before the commit is applied) and the version at $c$ itself (after the commit is applied). Comparing these two versions allows us to determine which quality issues the commit introduced or fixed. Our analysis focuses on source files written in Python, JavaScript, and TypeScript. We exclude files that are unlikely to reflect production code quality (e.g., tests, documentation, configuration files, auto-built artifacts, and vendored dependencies). Files are classified based on their paths and naming patterns (e.g., files under test/ or __tests__/ directories, or matching *_test.py). The full classification rules are included in our replication package.
Static Analysis.  For each AI-authored commit $c$ that modifies a source file $f$, we check out two versions of $f$ (i.e., the version at $c$’s parent commit (before) and the version after applying $c$ (after)). We run the same static analysis toolchain on both versions to identify potential code issues. We use ESLint (for JavaScript and TypeScript) [15] and Pylint (for Python) [52] to detect code smells and correctness issues. For security-related issues, we use Semgrep [61], which provides a unified framework for multi-language static analysis. For each detected issue, we record its rule identifier, line number, detector, and message. Figure 5 shows one such record, where ESLint flags a duplicate object key in a Claude-authored commit [64]. This produces two issue sets $I$: the issue set before the commit, denoted by $I_{f}^{-}$, and the issue set after the commit, denoted by $I_{f}^{+}$.
Repository: superagent-ai/superagent
Commit: 46695d1 (author: Claude)
File: node/src/server.js
Detector: ESLint
Rule: no-dupe-keys
Line: 86
Message: Duplicate key ‘lazyConnect’.
Figure 5: Example of a recorded issue detected by ESLint in a Claude-authored commit [64].